OpenStudy (anonymous):

simplify 3 ^3(sqrt) 24

OpenStudy (anonymous):

Please, try drawing it or use "Equation button" to inset you question.

OpenStudy (anonymous):

okay hold on

OpenStudy (anonymous):

Simplify\[3\sqrt[3]{24}\]

OpenStudy (anonymous):

\[3\sqrt[3]{24} = 3\sqrt[3]{8 * 3} = 2*3\sqrt[3]{3}= 6\sqrt[3]{3}\]

OpenStudy (anonymous):

Thnak you so much dude! :)

OpenStudy (anonymous):

You are welcome :)